OpenStudy (anonymous):

Solve by substitution: 2x-2y=-6 and y=x+3

OpenStudy (anonymous):

The second one states that y = x + 3 so you can plug that into the first equation. 2x - 2(x+3) = -6 2x - 2x - 6 = -6 Positive 2x and negative 2x cancel out, so you're left with -6 = -6

OpenStudy (anonymous):

Are there answer choices to the question or do you just have to work it out?